### v1

@@ A.1 Posts expressing Antidemocratic Attitudes and Partisan Animosity
We define AAPA based on eight variables identified [38] as potentially harmful to democratic functioning. While the previous study tested multiple interventions to measure the impact on these variables, we focus on content expressing support for these outcomes, following findings from an in-lab experiment [39].
@@
    Partisan animosity (v1): defined as dislike for opposing partisans [18].
@@
    Support for undemocratic practices (v2): defined as willingness to forgo democratic principles for partisan gain [61, 62]
@@
    Support for partisan violence (v3): defined as willingness to use violent tactics against outpartisans [63]
@@
    Support for undemocratic candidates (v4): defined as willingness to ignore democratic practices to elect inparty candidates [64]
@@
    Opposition to bipartisan cooperation (v5): defined as resistance to cross-partisan collaboration [65]
@@
    Social distrust (v6): defined as distrust of people in general [66]
@@
    Social distance (v7): defined as resistance to interpersonal contact with outpartisans [67]
@@
    Biased evaluation of politicized facts (v8): defined as skepticism of facts that favor the worldview of the other party [68]
@@
These definitions, evaluated in previous work [39], were used to prompt the LLM (see Table 2). A post is classified as AAPA when at least four of these factors are present. Over the course of 10 days, we found that, on average, 30.6% of the political posts in the pre-intervention feed of participants were classified as AAPA (Fig. 4). The distribution of AAPA varied by participants’ political orientation, with 34.25% of political posts in Republican feeds and 28.54% in Democratic feeds classified as AAPA (Fig. 5). The most common factors are Social Distance, Biased Evaluation of Politicized Facts, and Partisan Animosity (Fig. 6).
@@
Fig. 7 shows the correlation between these factors, showing that some of them have a high chance of occurring together. Fig. 8 shows that these factors cluster in 3 groups with {v2, v3, v4}, {v5}, and {v1, v6, v7, v8} tendency to appear together.
@@
Figure 4: Left: Distribution of the AAPA scores assigned to political posts in the “For you” feed on X (pre-intervention). The green bars represent the fraction of non-AAPA, while the orange bars describe the AAPA posts recommended by X.
Figure 5: AAPA score distribution for political posts, by party.
@@
Figure 6: Distribution of AAPA factors in political posts seen by participants.
@@
Figure 7: Correlation between the eight AAPA variables based on the co-occurrence in all the posts recommended by X to all participants in their ”For You” feed.
@@
Figure 8: PCA analysis of the eight AAPA variables. Reduction of the binary vectors representing the factors present in all the posts recommended by X to all participants in their ”For You” feed. The eight factors cluster into three groups.
A.2 Feed reranking implementation
@@
We developed a browser extension to run the field experiment intervening in real-time in the participants’ original feeds [47]. Our extension was distributed through the Google Chrome Store and supported Chrome and Edge browsers. Once the participants consented to participate, the extension is enabled every time they visit their feed on X. The extension is enabled only when the participants use the “For You” feed, which is the feed generated by the recommendation system. Participants can read their chronological feed at any time, but a tooltip regularly reminds them to switch to the “For You” tab to participate in the study. Our implementation relies on the browser extension to intercept the feed and on a remote backend to rescore and rerank the posts.
@@
Every time the user loads the feed, the extension receives a working set of 35 posts (5 of which are ads) from the server, and for each of these batches, the reranking procedure works as follows. First, the extension intercepts the response returned from the server and extracts the text of all the posts. If the post contains a link, we append the title and description visible in the preview with the prefix “Attached article’s description: ” and discard the URL. If the text quotes another post, we append the quoted text with the prefix “Quoting: .” Second, since our reranking logic is applied only to political posts, our reranking service selects all relevant posts in the feed using a RoBERTa-based political classifier (Sec. B.3). Filtering for political posts with a lighter model allows us to reduce the latency by scoring only the relevant posts. Then, the political posts are scored in real-time using GPT-3.5 using prompts based on previous work [39] and summarised in Table 2. For each post, the resulting data is a vector of 8 binary indicators flagging the presence of each factor. The final score of each post can vary between 0 and 8, and it is obtained by counting how many factors the classifier identified.
@@
Finally, depending on the participant’s exposure group, we apply the relevant intervention and return the modified feed to the browser, potentially with the in-feed survey. Each in-feed survey contained a slider for participants to respond using a 0–100 scale on the feeling thermometer. The design required two control conditions since the Reduced Exposure intervention placed the in-feed survey instrument where downranked content had previously been, and the Increased Exposure intervention placed the in-feed assessment instrument where upranked content would have been inserted.
A.2.1 Reduced Exposure feed
@@
Participants in the treatment condition of the Reduced Exposure group receive a feed with less AAPA content. This is achieved by penalizing these posts and moving them to a lower position in the feed. Once the relevant posts have been scored, we proceed as follows. First, we select one random AAPA item and flag the next position as the potential location for the in-feed survey. Then, if the participant is currently in the treatment condition, we downrank all the problematic posts. If the participant is in the control condition, we return the feed with the original ranking. The penalization offset for downranked posts depends on their score (position × score × 10) to ensure that posts with higher scores require more effort to become visible in the feed. Our implementation does not remove the posts, and the participants can read the downranked content by scrolling down the feed in a continuous session. This implementation ensures that the selection logic for the survey position is equivalent for the two experimental conditions, differing only in the downranking of the problematic posts for the treatment group. Finally, the in-feed survey is injected in the assigned position if the current feed load is selected for inserting the survey. If the current feed does not contain any problematic posts, there is no intervention, and we do not add any surveys. The effect of the intervention is visible in Fig. 9.
A.2.2 Increased Exposure feed
@@
The participants in the Increased Exposure group are presented with a treatment opposite to the Reduced Exposure group, as they can find AAPA content in the top positions of their feed. The treatment condition of this group receives a feed where AAPA posts are upranked. Since we cannot access the full inventory available on the platform to generate the custom recommendation, we source the AAPA posts from a custom inventory compiled from all the items recommended to the participants enrolled in the study. In particular, to ensure recency and relevance, we limit the candidate set to posts not already seen by the current participant and recommended by the feed algorithm to other participants with the same political leaning in the last 24 hours. This list may include posts recommended to the same participants that were not previously displayed. The content to uprank is randomly selected from the 100 posts with the highest AAPA score and matching these characteristics. Once the post is selected, we select one random position in the feed, and if the participant is in the treatment condition, the item will be inserted. Alternatively, if the participant is assigned to the control condition, the process is interrupted, and the post is not added. Despite not adding the post, mirroring the same procedure for the control group ensures the similarity of the two conditions in all respects, including potential server latency. Finally, if the event is sampled to include an in-feed survey, we append the in-feed survey in the next available position. The effect of the intervention is visible in Fig. 10.
Figure 9: Reduced Exposure: Distribution of the average number of political posts that each participant saw by day for each score. During the treatment period, participants experienced a significant reduction (-91.2%) in their exposure to AAPA content ($\geq$ 4 factors). While AAPA content was downranked, participants who scrolled far enough through their feeds still encountered a small amount of AAPA posts.
Figure 10: Increased Exposure: Distribution of the average number of political posts that each participant saw by day for each score. During the treatment period, participants experienced a significant increase (+23.1%) in their exposure to AAPA content ($\geq$ 4 factors).
Figure 11: Average AAPA score for political posts by day and condition; first averaged per participant and then averaged across all users in the corresponding group. The results demonstrate that the intervention increased and decreased AAPA exposure, as intended.
@@
Figure 12: Distribution of the number of AAPA posts upranked and downranked per participant, per day, by in the feeds of participants assigned to the corresponding treatment groups. The x-axis is on a logarithmic scale. The Reduce Exposure treatment downranked approximately an order of magnitude more AAPA posts per participant than the number of posts the Increased Exposure treatment upranked.
Appendix B Experimental platform
@@ B.1 Browser extension
The extension manages the ranking manipulation by modifying the content available on the X feed in real-time. Our implementation allows us to inject custom code on the main page to override the methods of the native XMLHttpRequest object. This approach allows us to intercept the communication with the server and modify the execution flow of the requests. Our browser extension subscribes to the requests endpoints used to retrieve the new content to add to the feed, interrupt the execution flow, send the server response to our backend for scoring and reranking, and finally restore the execution flow by calling the original methods with the response customized by our server. This approach requires careful understanding of the communication protocol with the server, but it offers significant advantages: first, the content can be modified by editing structured JSON objects without manipulating the final HTML, and second, by selecting what requests to intercept, it allows logging the interactions with the web page tracked by X without adding custom listeners.
@@
Finally, the extension manages the onboarding flow to ensure that participants can register and join the study only by following links that automatically guide them in installing the extension and setting the identifier from the recruitment platform [40].
B.2 Backend
@@
The backend is responsible for reranking the feed, logging the client events, sending daily participation reminders by email, and orchestrating the participants’ registration. The server is written in Python and Java using MongoDB and Memcached for storage and caching. When the browser extension interrupts the execution flow to request the reranking of the feed, X’s rotating circular loading indicator is extended until our reranking is complete. The implementation choices are guided by minimizing the response time to ensure minimal destruction of the user experience.
@@
The feed reranking proceeds in four steps: first, the response from X’s server, typically containing 35 posts, is sent to our backend and parsed to extract the text from the posts. The implementations focused on individual posts and in-feed conversations, ignoring the other types of content, such as ads or users’ recommendations. Second, all the posts are evaluated by a political classifier (Sec. B.3) to filter the posts that must be scored for the presence of the eight factors problematic for democratic functioning. This service runs as a standalone Python web service, accelerated by GPU, that introduces, on average, a latency of 490 ms (Std = 111 ms). Then, the political posts are dispatched to a Java server that manages an HTTP connection pool with OpenAI APIs. The service batches the posts to reduce the number of requests and uses a pool of tokens to minimize latency and avoid the rate limit exceptions. Finally, the scores are parsed and returned to the main backend, which applies the downranking intervention and adds the in-feed surveys. These steps are replicated in the same way in the control condition, but in that case, the backend returns the original feed. The downranked posts are saved in a cache (Memcached) and added to the feed when the user reaches the new position in a continuous scrolling session.
@@
The backend also stores the logs of interactions with the feed. When a participant interacts with the feed on X, the extension regularly shares with the backend what actions are performed on the main page. This data is limited to the actions on the main feed and does not include any private content. The actions recorded include when a post enters the main viewport, likes, favorites, reposts, clicks on external links, and the time spent on the browser tab.
@@
To facilitate communication with the participants in the registration flow, they are invited to share an email address. The platform uses this address to send automatic messages that guide the participant in completing the study successfully. These messages include a copy of the onboarding instructions and the consent form for future reference, a daily reminder sent at 6 PM (participant’s local time) inviting the participant to check their X feed, and the end-of-study notification with the link to the post-survey and the instruction to remove the browser extension.
@@
Finally, the backend orchestrates the registration flow of the participants [47]. Since the browser extension does not support passing custom installation parameters, the backend is responsible for guiding the onboarding flow to configure the extension for the current participant. After the termination of the pre-survey, the participants are redirected to our web server using a URL that encodes all the parameters relevant for the extension to work, such as the identifier and political position. Our server adds a browser cookie with this data and displays the step-by-step instructions for installing the extension. Participants are then redirected to the Chrome store, where they install the browser extension. As soon as the extension is installed, it opens a browser tab and loads to a welcome page hosted on our coordination server, where it has access to the cookies information previously set. This approach allows the extension to save these parameters in persistent storage without forcing the participants to enter this information manually after the installation of the extension. Lastly, the welcome message contains a link to redirect the participants back to the recruitment platform to set the task as completed.
B.3 Political Classifier
@@
Since our downranking intervention applies only to political posts, we developed a classifier to prefilter the relevant posts. Since GPT has a pay-per-use business model and calling an external API service call can introduce additional latency, we ensure that GPT is used to score only the content in which the eight variables are relevant. The political classifier is a distilled version of GPT-4 by fine-tuning a RoBERTa model to identify political posts, including mentions of officials and activists, social issues, news, and current events.
@@
The posts used to train the model were collected with the Twitter API in July 2023. To select the posts, we replicated the reverse chronological feeds of users across the political spectrum for two weeks and took a random sample of 50,000 posts. We first selected 300 users and then collected the posts by the users they followed.
@@
To select the users, we first obtained the followers of all members of Congress. Then, we categorized each user in the superset of followers based on how many Republican and Democrat members of Congress they followed: (a) far-left: followed 4-6 Democrats and no Republicans; (b) left: followed 1-3 Democrats and no Republicans; (c) middle: followed 1-3 Democrats and 1-3 Republicans; (d) far-middle: followed 4-6 Democrats and 4-6 Republicans; (e) right: follow no Democrats and 1-3 Republicans; (f) far-right: followed no Democrats and 4-6 Republicans. We started with members of Congress since their political affiliations are known and are often used to estimate Twitter users’ political leaning [69, 70]. Next, we removed follower accounts that are likely to be bots. We used the Botometer API and filtered out accounts with Complete Automation Probability above 0.6 [71]. Finally, we took a stratified sample of 300 accounts, selecting 50 users from each of the six groups defined above.
@@
We annotated all 50,000 posts by prompting GPT-4 using a definition proposed by the Pew Research Center [44]:
@@
We evaluate the model on a held-out dataset composed of 5,000 posts annotated by GPT-4 and a random subset of 300 posts annotated by humans. For each post, we collect three labels using Prolific and assign the value based on majority voting. The model evaluated on these testing sets achieves an F1 score of 93.6% on the GPT-4 labels and an F1 score of 91.3% on the human ground truth.
Ground truth Test set size Precision [95% CI] Recall [95% CI] F1 score
GPT-4 5,000 0.937 [0.931, 0.942] 0.920 [0.911, 0.928] 0.936
Human labels 300 0.921 [0.868, 0.966] 0.905 [0.847, 0.954] 0.913
Table 1: Performance of the fine-tuned RoBERTa political classifier. Bootstrapped 95% confidence intervals.
B.4 Scoring service
The scoring service is a Java backend responsible for returning the number of each political post’s eight problematic factors. It is designed to minimize latency. When a new request is received, it proceeds as follows: after the political classifier identifies the posts to score, they are sent to this service in one request as a list of textual items. This service first splits the list into chunks of at most ten items and then compiles eight prompts for each of them. The prompts are in the format:

### v2

@@ Acknowledgments
We define AAPA based on eight variables identified as potentially harmful to democratic functioning  (?). While the previous study tested multiple interventions to measure the impact on these variables, we focus on content expressing support for these outcomes, following findings from an in-lab experiment  (?).
@@
    Partisan animosity (v1): defined as dislike for opposing partisans  (?).
@@
    Support for undemocratic practices (v2): defined as willingness to forgo democratic principles for partisan gain  (?, ?)
@@
    Support for partisan violence (v3): defined as willingness to use violent tactics against outpartisans  (?)
@@
    Support for undemocratic candidates (v4): defined as willingness to ignore democratic practices to elect inparty candidates (?)
@@
    Opposition to bipartisan cooperation (v5): defined as resistance to cross-partisan collaboration (?)
@@
    Social distrust (v6): defined as distrust of people in general  (?)
@@
    Social distance (v7): defined as resistance to interpersonal contact with outpartisans  (?)
@@
    Biased evaluation of politicized facts (v8): defined as skepticism of facts that favor the worldview of the other party  (?)
Score Post Example
0 JUST IN: K. Harris has rolled out plans for a \$6,000 baby tax credit.
1 BREAKING — TRUMP CLAIMS he’ll sit for an interview with E. Musk this Monday night.
2 Kamala mocking Trump after he spirals in paranoia over her crowd numbers [fire emoji]. Just perfect
3 Drop a “MAGA” [US flag emoji] if you believe Trump will destroy Kamala H. in a Presidential debate
4 Trump literally boarding Epstein’s jet. #TrumpPedoFiles deserves to be trending again.
5 OMG!!! M. T. Greene just threw a tantrum at the GOP convention during an interview, yelling “you little liar.” “You’re to blame for President Trump nearly getting assassinated.
6 I don’t give a damn if Trump got hit or it was just broken glass. Who the fuck cares? He’s still a lying traitor, brain-dead scumbag, fraudulent asshole who stole the presidency. The fact that he STILL walks free is a goddamn disgrace and failure.
7 Fuck the NYT, fuck A. Schiff, fuck Rob R., fuck Clooney, fuck every rich, entitled bastard who thinks their whims outweigh 17M Democratic voters. Trump + #Project2025 will be on you.
8 FRAT HOE KAMALA: Ex-criminal court judge confirms Kamala Harris is a corrupt witch, corrupt her whole career, slept her way up, and isn’t Black. Biden’s a racist dog who once said “Negro kids are like bugs”. If Democrats win, we must respond
Table S1: Examples of Non-AAPA (Scores 0–3) and AAPA (Scores 4–8) posts. The posts were randomly selected from those containing the keywords Biden, Harris, Trump, Republican, or Democrat. To preserve the anonymity of the post authors, the example posts in this table have been modified from their original text to preserve the original authors’ intent as closely as possible while not being directly searchable on the web.
@@
These definitions, evaluated in previous work  (?), were used to prompt the LLM (Table S3). A post is classified as AAPA when at least four of these factors are present. Table S1 shows sample posts with different scores. Over the course of 10 days, on average, 30.6% of the political posts in the pre-intervention feed of participants were classified as AAPA (Fig. S2). The most common factors are Social Distance, Biased Evaluation of Politicized Facts, and Partisan Animosity (Fig. S4). The distribution of AAPA varied by participants’ political orientation, with 34.25% of political posts in Republican feeds and 28.54% in Democratic feeds classified as AAPA (Fig. S3). These percentages vary over time and can be influenced by major political events Fig. S7. For instance, during the experiment, external events (such as the presidential nomination of Kamala Harris) were followed by an increase in the proportion of AAPA content, particularly in the feeds of Republican participants.
@@
Fig. S5 shows the correlation between these factors, showing that some of them have a high chance of occurring together. Fig. S6 shows that these factors cluster in 3 groups: {v2, v3, v4}, {v5}, and {v1, v6, v7, v8}, suggesting that they tend to appear together.
@@
Figure S2: Distribution of the AAPA scores assigned to political posts in the “For you” feed on X (pre-intervention). The green bars represent the fraction of non-AAPA, while the orange bars describe the AAPA posts recommended by X (not necessarily seen by the users).
Figure S3: AAPA score distribution for political posts, by party.
@@
Figure S4: Distribution of AAPA factors in political posts seen by participants.
@@
Figure S5: Correlation between the eight AAPA variables based on the co-occurrence in all the posts recommended by X to all participants in their “For you” feed.
@@
Figure S6: PCA analysis of the eight AAPA variables. Reduction of the binary vectors representing the factors present in all the posts recommended by X to all participants in their “For you” feed. The eight factors cluster into three groups.
Figure S7: Average fraction of AAPA content seen over time among political posts for participants in the two control groups (including during the baseline period), separated by party. Major political events are annotated. Shaded bands represent 95% bootstrapped confidence intervals.
S3.2 Feed Reranking Implementation
@@
We developed a browser extension to run the field experiment intervening in real time in the participants’ original feeds  (?). Our extension was distributed through the Google Chrome Store and supported Chrome and Edge browsers. Once the participants consented to participate, the extension is enabled every time they visit their feed on X. The extension is enabled only when the participants use the “For you” feed, which is the feed generated by the recommendation system. Participants can read their chronological feed at any time, but a tooltip regularly reminds them to switch to the “For you” tab to participate in the study. Our implementation relies on the browser extension to intercept the feed and on a remote backend to rescore and rerank the posts.
@@
Every time the user loads the feed, the extension receives a working set of 35 posts (5 of which are ads) from the server, and for each of these batches, the reranking procedure works as follows. First, the extension intercepts the response returned from the server and extracts the text of all the posts. If the post contains a link, we append the title and description visible in the preview with the prefix “Attached article’s description: ” and discard the URL. If the text quotes another post, we append the quoted text with the prefix “Quoting: .” Second, since our reranking logic is applied only to political posts, our reranking service selects all relevant posts in the feed using a RoBERTa-based political classifier (Appendix S4.3). Filtering for political posts with a lighter model allows us to reduce the latency by scoring only the relevant posts. Then, the political posts are scored in real-time using GPT-3.5 using prompts based on previous work  (?) and summarized in Table S3. For each post, the resulting data is a vector of 8 binary indicators flagging the presence of each factor. The final score of each post can vary between 0 and 8, and it is obtained by counting how many factors the classifier identified.
@@
Finally, depending on the participant’s exposure group, we apply the relevant intervention and return the modified feed to the browser, potentially with an in-feed survey. Each in-feed survey contained a slider for participants to respond using a 0–100 scale on the feeling thermometer. The design required two control conditions since the Reduced Exposure intervention placed the in-feed survey instrument where down-ranked content had previously been, and the Increased Exposure intervention placed the in-feed assessment instrument where up-ranked content would have been inserted.
S3.2.1 Reduced Exposure Feed
@@
Participants in the treatment condition of the Reduced Exposure group receive a feed with less AAPA content. This is achieved by penalizing these posts and moving them to a lower position in the feed. Once the relevant posts have been scored, we proceed as follows. First, we select one random AAPA item and flag the next position as the potential location for the in-feed survey. Then, if the participant is currently in the treatment condition, we down-rank all AAPA posts. If the participant is in the control condition, we return the feed with the original ranking. The penalization offset for down-ranked posts depends on their score (position × score × 10) to ensure that posts with higher scores require more effort to be viewed in the feed. Our implementation does not remove the posts, and the participants can read the down-ranked content by scrolling down the feed in a continuous session. This implementation ensures that the selection logic for the survey position is equivalent for the two experimental conditions, differing only in the down-ranking of AAPA posts for the treatment group. Finally, the in-feed survey is injected in the assigned position if the current feed load is selected for inserting the survey. If the current feed does not contain any AAPA posts, there is no intervention, and we do not add any surveys.
S3.2.2 Increased Exposure Feed
@@
The participants in the Increased Exposure group are presented with a treatment opposite to the Reduced Exposure group, as they can find AAPA content in the top positions of their feed. The treatment condition of this group receives a feed where AAPA posts are up-ranked. Since we cannot access the full inventory available on the platform to generate the custom recommendation, we source the AAPA posts from a custom inventory compiled from all the items recommended to the participants enrolled in the study. In particular, to ensure recency and relevance, we limit the candidate set to posts not already seen by the current participant and recommended by the feed algorithm to other participants with the same political leaning in the last 24 hours. This list may include posts recommended to the same participants that were not previously displayed. The content to up-rank is randomly selected from the 100 posts with the highest AAPA score matching these characteristics. Once the post is selected, we select one random position in the feed, and if the participant is in the treatment condition, the item is inserted there. Alternatively, if the participant is assigned to the control condition, the post is not added. Despite not adding the post, mirroring the same procedure for the control group ensures the similarity of the two conditions in all respects, including potential server latency. Finally, if the event is sampled to include an in-feed survey, we append the in-feed survey in the next available position.
S3.2.3 Manipulation Checks
To assess whether the interventions altered participants’ feeds as intended, we analyzed both the number of AAPA posts that were up-ranked or down-ranked and the resulting exposure to political and AAPA content.
To avoid excessive exposure to potentially polarizing content, the two experiments were not symmetric in terms of the amount of content reranked: in the Reduced Exposure condition, all AAPA posts detected in a participant’s feed were down-ranked, while in the Increased Exposure condition, only one high-scoring AAPA post was up-ranked per batch (typically containing 30 posts). As shown in Fig. S8, this design resulted in a substantial difference in the number of reranked posts per user per day, with the Reduced Exposure condition reranking significantly more content. Because AAPA content was already prevalent during the study period, reducing exposure was achieved by down-ranking many posts, whereas increasing exposure required up-ranking fewer but more intense AAPA posts, making this condition a comparatively harder test.
Fig. S9 shows the effect of the interventions on overall feed composition, summarizing the fraction of AAPA posts in participants’ feeds for each day of the experiment. During the three-day baseline period, participants in all four experimental groups were exposed to a similar fraction of AAPA posts per day. However, during the subsequent seven-day intervention period, participants in the two control groups continued to be exposed to a similar fraction of AAPA posts, whereas participants in the treatment groups of the Reduced and Increased Exposure experiments experienced significant decreases and increases in AAPA exposure, respectively. As discussed above, the Increased Exposure treatment up-ranked fewer AAPA posts, and thus the relative increase in AAPA exposure is smaller than the decrease observed in the Reduced Exposure treatment. We note that the fraction of AAPA posts in the Reduced Exposure treatment does not fall all the way to zero, since posts are down-ranked rather than removed, and participants can still reach them by scrolling much further down their feeds. Importantly, the interventions specifically targeted AAPA content: as shown in Fig. S10, participants continued to be exposed to substantial amounts of political content in both experiments. In the Reduced Exposure treatment, the decline in exposure to political content is limited to the subset of AAPA content, whereas in the Increased Exposure condition, the overall volume of political content remains nearly unchanged.
Figure S11 shows the average AAPA score of all political posts to which participants were exposed to in each of the four experimental groups. Although the Increased Exposure treatment up-ranked fewer AAPA posts, it focused on posts with high AAPA scores. As a result, the Reduced and Increased Exposure treatments induced a similar relative change in average AAPA scores in the respective directions.
Finally, Fig. S12 and Fig. S13 show the daily average exposure to political posts by AAPA score, separately for treatment and control groups in each experiment. In the Reduced Exposure treatment, we observe a clear drop in the number of high-AAPA posts seen by participants after the intervention begins, while the control group remains stable. In contrast, the Increased Exposure treatment shows a rise in exposure to high-AAPA posts at the beginning of the intervention period.
Overall, these results suggest that the treatments successfully induced the intended variation in AAPA exposure.
@@
Figure S8: Distribution of the number of AAPA posts up-ranked and down-ranked per participant, per day, in the feeds of participants assigned to the corresponding treatment groups. The horizontal axis is on a logarithmic scale. The Reduce Exposure treatment down-ranked approximately an order of magnitude more AAPA posts per participant than the number of posts the Increased Exposure treatment up-ranked.
Figure S9: Average fraction of AAPA posts seen by participants for each day of the experiment. In the Reduced Exposure condition, the fraction drops at the start of the treatment period but does not reach zero, as some down-ranked posts still appear with sufficient scrolling.
Figure S10: Average fraction of political posts seen by participants for each day of the experiment. Despite some variation, participants’ feeds in both experiments continued to contain a significant amount of political content.
Figure S11: Mean AAPA score for political posts by day and condition; first averaged per participant and then averaged across all users in the corresponding group. The results demonstrate that the intervention increased and decreased AAPA exposure, as intended.
Figure S12: Reduced Exposure: Distribution of the average number of political posts that each participant saw by day for each score. During the treatment period, participants experienced a significant reduction (-89.6%) in their exposure to AAPA content ($\geq$ 4 factors). While AAPA content was down-ranked, participants who scrolled far enough through their feeds still encountered a small amount of AAPA posts.
Figure S13: Increased Exposure: Distribution of the average number of political posts that each participant saw by day for each score. During the treatment period, participants experienced a significant increase (+26.4%) in their exposure to AAPA content ($\geq$ 4 factors).
S4 Experimental Platform
@@
The extension manages the ranking manipulation by modifying the content available on the X feed in real time. Our implementation allows us to inject custom code on the main page to override the methods of the native XMLHttpRequest object. This approach allows us to intercept the communication with the server and modify the execution flow of the requests. Our browser extension subscribes to the request endpoints used to retrieve the new content to add to the feed, interrupts the execution flow, sends the server response to our backend for scoring and reranking, and finally restores the execution flow by calling the original methods with the response customized by our server. This approach requires careful understanding of the communication protocol with the server, but it offers significant advantages: first, the content can be modified by editing structured JSON objects without manipulating the final HTML, and second, by selecting what requests to intercept, it allows logging the interactions with the web page tracked by X without adding custom listeners.
@@
Finally, the extension manages the onboarding flow to ensure that participants can register and join the study only by following links that automatically guide them in installing the extension and setting the identifier from the recruitment platform  (?).
In this work, we have focused on desktop-based usage, but a similar approach can be implemented on mobile phones through web views and emerging support for mobile browser extensions (e.g., Microsoft Edge for Mobile).
S4.2 Backend
@@
The backend is responsible for reranking the feed, logging the client events, sending daily participation reminders by email, and orchestrating the participants’ registration. The server is written in Python and Java using MongoDB and Memcached for storage and caching. When the browser extension interrupts the execution flow to request the reranking of the feed, X’s rotating loading indicator is extended until our reranking is complete. The implementation choices are guided by minimizing response time to ensure minimal disruption to the user experience.
@@
The feed reranking proceeds in four steps: first, the response from X’s server, typically containing 35 posts, is sent to our backend and parsed to extract the text from the posts. The implementation focused on individual posts and in-feed conversations, ignoring the other types of content, such as ads or users’ recommendations. Second, all the posts are evaluated by a political classifier (Appendix S4.3) to filter the posts that need to be scored for the presence of the eight AAPA factors. This service runs as a standalone Python web service, accelerated by a GPU, that introduces, on average, a latency of 490 ms (STD = 111 ms). Then, the political posts are dispatched to a Java server that manages an HTTP connection pool with OpenAI APIs. The service batches multiple posts in a single prompt to reduce the number of requests and uses a pool of tokens to minimize latency and manage rate limits. Finally, the scores are parsed and returned to the main backend, which applies the reranking intervention and adds the in-feed surveys. These steps are replicated in the same way in the control condition, but in that case, the backend returns the original feed. In the case of down-ranking, the posts are saved in a cache (Memcached) and added to the feed when the user reaches the new position in a continuous scrolling session.
@@
The backend also stores the logs of interactions with the feed. When a participant interacts with the feed on X, the extension regularly shares with the backend what actions are performed on the main page. This data is limited to the actions on the main feed and does not include any private content. The actions recorded include when a post enters the main viewport, likes, reposts, clicks on external links, and the time spent on the browser tab.
@@
To facilitate communication with the participants in the registration flow, they are invited to share an email address. The platform uses this address to send automatic messages that guide the participant in completing the study successfully. These messages include a copy of the onboarding instructions and the consent form for future reference, a daily reminder that they are enrolled in the study sent at 6 PM (participant’s local time), and the end-of-study notification with the link to the post-experiment survey and the instruction to remove the browser extension.
@@
Finally, the backend orchestrates the registration flow of the participants  (?). Since the browser extension does not support passing custom installation parameters, the backend is responsible for guiding the onboarding flow to configure the extension for the current participant. After the completion of the pre-experiment survey, the participants are redirected to our web server using a URL that encodes all the parameters relevant for the extension to work, such as the identifier and political position. Our server adds a browser cookie with this data and displays the step-by-step instructions for installing the extension. Participants are then redirected to the Chrome store, where they install the browser extension. As soon as the extension is installed, it opens a browser tab and loads to a welcome page hosted on our coordination server, where it has access to the cookies information previously set. This approach allows the extension to save these parameters in persistent storage without forcing the participants to enter this information manually after the installation of the extension. Lastly, the welcome message contains a link to redirect the participants back to the recruitment platform to set the task as completed.
S4.3 Political Classifier
@@
Since our down-ranking intervention applies only to political posts, we developed a classifier to prefilter the relevant posts. Since GPT has a pay-per-use business model and calling an external API service can introduce additional latency, we ensure that GPT is used to score only the content in which the eight variables are relevant. The political classifier is a distilled language model fine-tuning a RoBERTa model to identify political posts, including mentions of officials and activists, social issues, news, and current events, based on an annotated training set.
@@
The posts used to train the model were collected with the X API in July 2023. To select the posts, we replicated the reverse chronological feeds of users across the political spectrum for two weeks and took a random sample of 50,000 posts. We first selected 300 users and then collected the posts by the users they followed.
@@
To select the users, we first obtained the followers of all members of Congress. Then, we categorized each user in the superset of followers based on how many Republican and Democrat members of Congress they followed: (a) far-left: followed 4-6 Democrats and no Republicans; (b) left: followed 1-3 Democrats and no Republicans; (c) middle: followed 1-3 Democrats and 1-3 Republicans; (d) far-middle: followed 4-6 Democrats and 4-6 Republicans; (e) right: followed no Democrats and 1-3 Republicans; (f) far-right: followed no Democrats and 4-6 Republicans. We started with members of Congress since their political affiliations are known and are often used to estimate the political leanings of users on X  (?, ?). Next, we removed follower accounts that are likely to be bots. We used the Botometer API and filtered out accounts with Complete Automation Probability above 0.6  (?). Finally, we took a stratified sample of 300 accounts, selecting 50 users from each of the six groups defined above.
@@
Ground truth Test set size Precision [95% CI] Recall [95% CI] F1 score
GPT-4 5,000 0.937 [0.931, 0.942] 0.920 [0.911, 0.928] 0.936
Human labels 300 0.921 [0.868, 0.966] 0.905 [0.847, 0.954] 0.913
Table S2: Performance of the fine-tuned RoBERTa political classifier. Bootstrapped 95% confidence intervals.
We annotated all 50,000 posts by prompting GPT-4 using a definition proposed by the Pew Research Center  (?):
@@
We evaluate the model on a held-out dataset composed of 5,000 posts annotated by GPT-4 and a random subset of 300 posts annotated by humans. For each post, we collect three labels using Prolific and assign the value based on majority voting. The model achieves an F1 score of 93.6% on the GPT-4 labels and an F1 score of 91.3% on the human ground truth.
S4.4 Scoring Service
The scoring service is a Java backend responsible for evaluating each political post across the eight AAPA dimensions. It is designed to minimize latency. When a new request is received, it proceeds as follows: after the political classifier identifies the posts to score, they are sent to this service in one request as a list of textual items. This service first splits the list into chunks of at most ten items and then compiles eight prompts for each of them. The prompts are in the format: